OpenStudy (anonymous):

Stacie paid $13.25 for new earphones. The price she paid included a sales tax of 6%. What was the price of the earphones before the tax was added? A. $12.46 B. $12.50 C. $13.19 D. $14.05 OH PLEASE HELP ME! I WILL GIVE MEDAL

OpenStudy (anonymous):

@AdamK @akonkel @AdavisKTMDamTheBam

OpenStudy (anonymous):

plz help

OpenStudy (adamk):

13.25 x 0.06 Then, subtract that amount from 13.25

OpenStudy (ribhu):

12.5

OpenStudy (adamk):

@ribhu Please do not give direct answers.

OpenStudy (anonymous):

thanks

OpenStudy (ribhu):

wron method @adam is telling

OpenStudy (anonymous):

but the answer is correct

OpenStudy (anonymous):

just the method he did was wrong

OpenStudy (anonymous):

oh but @AdamK is really good

OpenStudy (ribhu):

yeah and the previous one also was wrong

OpenStudy (anonymous):

he is the best at helping people, especially me

OpenStudy (anonymous):

okay

OpenStudy (ribhu):

13.25/(1+.06)

OpenStudy (anonymous):

the answer u mean

OpenStudy (adamk):

I'm not wrong. I did it correctly.

OpenStudy (ribhu):

see let the price be x, tax is 6% so 1.06x = 13.25 hence x = 13.25/1.06

OpenStudy (adamk):

13.25 x 0.06 = 0.795 13.25 - 0.795 = 12.455 which rounds to 12.50

OpenStudy (adamk):

There are multiple ways how to do it.

OpenStudy (ribhu):

you check you told 13.25*(1-.06) = 12.455

OpenStudy (anonymous):

i did what u did @AdamK

OpenStudy (anonymous):

just like that

OpenStudy (ribhu):

but it is wrong just think what r u doing? tax is calculated on the price and you are calculating the tax on already tax included price. @AdamK

OpenStudy (ribhu):

just read it well and then blame. @Star222 @AdamK

OpenStudy (ribhu):

this doesn't require rounding but there comes exact answer 12.5

OpenStudy (ribhu):

@AdamK with your method answer comes to be A( 12.555) and nt B(12.50)